Grade 10Physical Chemistry

Calculate molality and molarity of h2so4 containing 27% w/w and its density is 1.2gm/ml

Hansa Kachhatiya
9 Years agoGrade 10

2 Answers

Vaibhav Singh
9 Years ago
27℅w/w H2SO4 means 27 g of H2SO4in 100 g of solution. Therefore moles of h2so4= 27/98 =0.27. density of solution= 1.2g/ml.1.2 = 100/Vtherefore V = 100/1.2 ml or 1/12 L=0.083LNow molarity of h2so4 = 0.27/0.083= 3.25 Mand molality = 0.27/0.073 = 3.69m
Sachin Kewalramani
9 Years ago
Hi,Molarity = %W/W * 10 * D/Molar Mass of soluteTherefore, Molarity = 27*10*1.2/98 =3.306 MMolality = 1000 * Molarity / (1000d-M) * Molar mass of soluteTherefore, Molality = 3306/(1200-3.306) * 98 =0.3773 m